OpenStudy (anonymous):

Let f^-1(x)= 1/2x-5 Which points lie on the graph of y=f(x)? Show all your work A (-4,-7),(6,-2) B (-6,-2), (1,12) C (-1,-12), (2, -6) D (-8, 1), (2,6)

OpenStudy (anonymous):

\[f^-1(x)= \frac{ 1 }{ 2 }x-5\]

OpenStudy (anonymous):

the answer is B \[f^-1(x)=\frac{ 1 }{ 2 }x-5\] let f^-1(x)=y \[y=\frac{ 1 }{ 2 }x-5\] \[y+5=\frac{ 1 }{ 2 }x\] 2(y+5)=x so, f(x)=2(x+5) f(x)=2x+5 Substituting x=-6, f(-6)=2(-6)+10 f(-6)=-2 Substituting x=1 f(1)=2(1)+10 f(1)=12 So, the graph lies on the line (-6,-2) and (1,12)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome :)